Wednesday, November 30, 2011

Carlo's Scribe Post- Textbook 4,5 and 8

4)For 4 A what I did was I divided 20 and 100 then I got 0.2After that I multiply 0.2 to 60 so my final answer is 12.





B) For 4 B what I did was the same thing as 4 A. I divided 250 to 100 and I get 2.4 and then I multiply 2.4 to 400 so my final answer is 1000

.























C)
For 4 C I did it the different way what I did was a proportion statement so what I did was I wrote 10.50 and divided it by 100 and get 0.105 same thing to the bottom I divided it by 100. After I got the answer I took 0.105 and multiply it by 100 and get the final answer of 10.5.















5) For 5 A what I did was I took 2,5 and divided it together so then I get 0.4 then after I divided 0.4 to 100 and get 0.004, after I multiply 0.004 to 32.5 and the answer of 1.3.












B) For 5 B what I did was the same thing with 5A but I didn't move any decimals.












C)For 5 C is pretty much the same thing as what I did earlier. You divide 175 to 100 and get 1.75 then after that I multiplied 1.75 to 25.50 and get $219.625.

8) What I did for 8 was a ratio table, First thing I did was 100 divide by 4 then I get 25, 100 divide by 100 then I get 1 and the last what I divide 1 by 2 and get 1/2. The same thing to the other side what ever you do to the one side, you do the same thing on the other side so my final answer is $21.42.















Comment If I did something wrong! Thank you (:

John's Scribe Post-Textbook 10, 13 and 14

10) When water freezes, its volume increases
by approximately 10%.
a) By how much does the volume of a) 10
-- = 0.1 0.1x750=75mL
750 mL of water increase when it 100
freezes?

b) What is the volume of ice created? b) 825mL


14) 4% of 100 is the same as 8% of what
number? Explain how you arrived at
your answer.

I arrive at 4% is half of 8% and 50 is half of 100

13) Suppose a real estate agent receives 5%
commission on the fi rst $200 000 of a
house’s selling price, and 6% on the
remaining amount.
a) What does commission mean? a) Commission is the portion of the sale price that the real estate agent earns.

b) If a house sells for $345 000, how
much commission does the real estate
agent make on the sale of the house? b) $18 700

Comment if I did something wrong.

Sarissa's Scribe Post - Textbook questions 3,7&11

Pages 142-143
Questions 3,7 & 11



a) 300/100 =3 b) 1.25/100=0.0125
3x 2000 = 6000 0.0125x 60= 0.75

c) 1/100=0.001
0.001x 40= 0.4




a)If there were 100 tickets there would be a 1% chance of winning with one ticket. So, 1/2 = 0.5 or 0.5%

b)If there is a 0.5% chance of winning with one ticket you would have to buy 5 tickets to have a 2.5% chance.
(0.5x5=2.5)







6.5 x 9 984 670 = 64 900 355 km2

JM's Scribe Textbook 4.3 # 6,10,14

Page 142 - 143
Numbers 6,10,14










6.A


6.B

6.C


10. A
10.B
I'm not sure how to do this so I've tried
Sorry =(


4% of 100
8% of 200

I arrived at my answer because 4% x 2 = 8%
100 x 2 = 200

If I did something wrong help me out, comment. =)


Ayra's Scribe Post 5, 9, 11









* 5. Determine the percent of each number.


Give your answer to the nearest hundredth.*





a) 1.3 b) approximately 144.88 c) $219.63

___________________________________________________________________


9. The highest point in Canada is Mount Logan, which is in the Yukon Territory. Mount Logan is 159% as high as the highest point in Alberta, Mount Columbia. The elevation of Mount Columbia is 3747 m. What is the elevation of Mount Logan?


answer - 5957.73 m





















____________________________________________________________________



11. The area of Canada is approximately 9 984 670 km2. The area of Manitoba is about 6 1/2% of the area of Canada.



What is the area of Manitoba to the nearest square kilometre?



answer - approximately 649 004 km2.

____________________________________________________________________



Here is a video: http://www.youtube.com/watch?v=2rr-IXInEVc Enjoy!

*Sorry if the fractions don't look like fractions.*

Tuesday, November 29, 2011

Kate's Pay it Forward

Part 1

Pay it forward means doing good things to other people and making a difference for their life.

Part 2

My Pay It Forward is an act of kindness to help Charry’s auntie's friend who's disabled. we did this with my partner Charry Mangaya. We decided what are we going to do. We chose this activity because we like to help (old) people with disabilities. We helped Precilla Malla, we call her tita (auntie) Malla, she's 59 years old, and she’s disabled with psoriasis arthritis and diabetes since 2003. We asked her what we can do for her. We did this act of kindness last November 13, 2011 at 11:30 A.M to 3:30 P.M.

Part 3

Our act of kindness went well. First Charry introduced me to her; we asked her "what can we do for her?”

We saw her kitchen, there's a lots of dirty dishes. So, we washed it, we cooked rice for her, we swept the floor, we cleaned her table. We saw her bedroom and she asked us is it okay to help her fold some of her clothes and we said “yes we can". We noticed that her bedroom is crowded with her things and a little bit messy; I told her that we can do some "interior designing to her bed room ". We moved some stuff, while we're cleaning up; we were talking and sharing stories like her experience when she's young. We're done at that time. We asked permission, if we can take a video and a photo to her. She approved. She was so happy when she saw her bedroom was all cleaned up. She told us that she's going to pay but we told her that it’s okay because we did this for free and she doesn't have to pay us .She's so thankful to us . She wants us to visit and to come back if we have a free time. Charry and I had so much fun . were so happy because we make her so happy and Charry and I helped each other to take pictures and videos and to help Tita Malla.

Part 4

Pay it forward is important because you share good things to others with the intension to make a difference not only to their life’s but to yourself as well .



Here's the link for our Pay It Forward Video.

http://www.youtube.com/watch?v=tVqDah52PN8

Gielh's home work book questions Scribe post




5) The commission for the sale of a house was 6 and 3 fourths %. IF the house sold for $345 000 how much was the commission?


How I got the answer: I divided 6.75 by 100 then i got the decimal number which
is 0.0675. I multiplied 0.0675 by 345000 then i got 23287. The commission is $23287.




7 a) Miguel Bought a car for $4700. He made a down payment of 19 1/2%. How much was the down payment
How I got the answer : I divided 19.5 by 100 so I got a decimal number which is 0.195. I multiplied 0.195 by 4700 and I got 916.5. So Miguel's down payment was $ 916.50.

b) about 5.6% of Canadians have type 2 diabetes . IF Canada's population is 32 million , About ho
w many Canadians have this Condition?

How I got the answers: I divide 5.6 by 100 so then I got a decimal number which is 0.056. I multiplied 0.056 by 32 million and then I got 1.792 million Canadians have diabetes.

c) the 4900 - seat hockey arena was 63% full. How many people were at the game?

How I got the answer : I divide 63 by a 100 then I got a decimal number which is o.63 % . I multiplied 0.63 by 4900 and i got 3087. So 3087 people .

8)The Nile river is about 209% the length of the Yukon River. If the Yukon River is 3168 km , How long is the Nile River ( to the nearest km?)

How I got the Answer : I Divide 209 by 100 so then I got a decimal number 2.09. Then I multiplied 2.9 by 3168 which is 6621.12 %

*I divided by 100 first because a percent is out of a 100.

Here is a video of how to find a percent of a number :


Hey guys I'm sorry If my writing was bad .